OpenStudy (anonymous):

Each time you click a toggle switch, the switch either turns from o ff to on or from on to off . Suppose that you start with three toggle switches with one of them on and two of them off . On each move you randomly select one of the three switches and click it. Let m and n be relatively prime positive integers so that m/n is the probability that after four such clicks, one switch will be on and two of them will be off . Find m + n.

OpenStudy (anonymous):

help plz?? anyone?

OpenStudy (anonymous):

um is the 3 switch

OpenStudy (anonymous):

Hold on I think I can get you an answer.

OpenStudy (anonymous):

lol its 3d switch

OpenStudy (anonymous):

There are 3 options to choose from so start with 3 There are 4 times you may try this Times 3 four times 3*3*3*3 This equals 27 With only one possible way to get it to one on two off, you are left with 1/27

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

this is not the right answer..

OpenStudy (paxpolaris):

\[{2+2+1+2+1 \over 5+4+3+2+1}= {8 \over 15}\] m+n= 8+15 =23 how about 23?

OpenStudy (anonymous):

neither.. :S

OpenStudy (paxpolaris):

sorry, this should be right: let the total number of outcomes be 3^4=81 eg. with 3 switches X,B,C we would have XXXX, XXXB, XBXX, XCXX, BCCB... etc \[\frac mn ={10+32+12+8+1 \over 16 +32+24+8+1}= \frac{63}{81}=\Large\frac 79\] m+n=15

OpenStudy (paxpolaris):

**m+n=16